# FedCQA: Answering Complex Queries on Multi-Source Knowledge Graphs via Federated Learning

### v2

@@ 5.1.3. Baselines
Figure 4. The evaluation results of FedCQA-GQE facing different types of cross-graph queries on FB15k-237 (subfigures (a), (b)). The evaluation results of FedCQA-GQE model facing in-graph queries on FB15k-237 (subfigures (c), (d)).
5.1.4. Evaluation Metrics
Following the previous work (Bai et al., 2023b), we evaluate the generalization capability of models by calculating the rankings of answers that cannot be directly retrieved from an observed knowledge graph. Given a testing query $q$, the training, validation, and public testing answers are denoted as $M_{train}$, $M_{val}$, and $M_{test}$, respectively. We evaluate the quality of retrieved answers using Hit ratio (HR) and Mean reciprocal rank (MRR). HR@K metric evaluates the accuracy of retrieval by measuring the percentage of correct hits among the top K retrieved items. The MRR metric evaluates the performance of a ranking model by computing the average reciprocal rank of the first relevant item in a ranked list of results. The metric can be defined as:
$m(r)=\textbf{1}[r\leq K]$ if the metric is HR@K and $m(r)=\frac{1}{r}$ if the metric is MRR. Higher values denote better reasoning performance. We train local models at each client by using the in-graph training queries and tune hyper-parameters using the validation queries. The evaluation is then finally conducted on the testing queries, including the evaluation of in-graph queries on local query encoding models and cross-graph queries on the global federated knowledge graph query answering system, respectively.
@@ 5.2. Performance Evaluation
We evaluate FedCQA’s complex query answering performance on three datasets and compare it to other baselines. We apply FedCQA on three base CQA models and evaluate the average retrieval performance on various queries. The results are summarized in Table 3 and Table 4. Table 3 reports the retrieval performance of various distributed knowledge graph complex query answering models when there are 3 clients. For each model, we evaluate performance facing in-graph queries and cross-graph queries, respectively. For in-graph queries, the average scores of all clients are reported. We report results in HR@3 and MRR which higher scores indicate better performance. The best results are underlined. The best results of distributed models are in bold. As shown in Table 3, our proposed methods can effectively retrieve complex query answers from distributed knowledge graphs. In comparison to local settings, we can see that FedCQA can utilize all participated knowledge graphs and performs better facing in-graph queries. For example, GQE model with FedCQA can retrieve 14.21 HR@3 on average while can only reach 11.85 without collaboration. Besides, in comparison to other federated knowledge graph embedding methods, our proposed FedCQA can reach comparable performance in both in-graph queries and cross-graph queries without exposing sensitive entity embeddings to the server. For example, FedCQA achieves the best performance in cross-graph queries in more than half of datasets and base CQA models.
In Table 4, we present the performance of FedCQA when there are 5 clients. We compare the performance with local training without collaboration to demonstrate the influence of client numbers. As shown in the table, FedCQA performs well compared to complex query answering models without collaboration, there are performance improvements in all datasets after applying FedCQA to various base CQA models, demonstrating that FedCQA can utilize the intrinsic information in the distributed knowledge sets. The collaboration in distributed knowledge graphs allows FedCQA to reason over various logical paths to improve performance.
@@ 5.3. Query Types
FedCQA can globally reason over distributed knowledge graphs and retrieve answers to cross-graph queries. To evaluate FedCQA’s performance on various types of complex queries, we conduct experiments to evaluate the retrieval performance of FedCQA and compare it to central learning on FB15k-237 when there are 3 clients. Figure 4(a) and Figure 4(b) show the FedCQA-GQE’s performance facing cross-graph queries. As we can see from the figure, FedCQA performs well on most various types of queries compared to the central model. For example, on query types ’2i’ and ’pi’, FedCQA can reach more than 90% MRR compared to the central model.
@@ 5.4. Local Influence
Because the in-graph queries can be processed by a single local knowledge graph CQA model, in this part, we evaluate the performance of FedCQA on these queries to assess the influence of FedCQA on local queries. We conduct experiments on FB15k-237 and the number of clients is 5. We evaluate the performance of FedCQA based on GQE and compare the model with no collaboration. The results are summarized in the Figure 4(c) and Figure 4(d). As shown in the figure, although each client has different performance due to the triplets correlation being different in each sub-dataset, FedCQA can improve all clients’ performance compared to the model without collaboration.
@@ 6. Conclusion
In this work, we present a federated complex query answering framework, FedCQA to reason over distributed knowledge sets with privacy preserved. We define the distributed knowledge graph complex query answering problem. To solve the problem, we propose secret aggregation for federated learning where the aggregated parameters can be kept secret to the server. Besides, we design a distributed query retrieval process for answering queries from distributed knowledge graph sets to protect clients’ privacy. To evaluate FedCQA model performance, we construct a benchmark based on three commonly used knowledge graph complex query answering datasets: FB15k-237, FB15k, and NELL. Extensive experiments on the benchmark demonstrate the effectiveness of our proposed FedCQA. FedCQA can retrieve answers given a query while keeping the sensitive information secret at local knowledge graphs. In the future, we aim to propose new methods for better answering complex queries by exploiting intrinsic information in the distributed knowledge graphs.
ACKNOWLEDGMENTS
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20) and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
@@ Appendix A Parameter Sharing
  • Client B chooses a secret integer $b$, then sends Alice $n_{B}=g^{b}\mod p$.
@@
Appendix B Algorithm
In this section, we present the pseudo code for FedCQA.
@@ Appendix B Algorithm
selected in each round $F$;
@@
  Upload $\mathbf{E}^{i}$ to server for secret aggregation, receive $\mathbf{E}_{0}^{i}$
@@
Algorithm 2 FedCQA Framework

### v3

@@ 5.1.4. Evaluation Metrics
Figure 4. The evaluation results of FedNGDB-GQE facing different types of cross-graph queries on FB15k-237 (subfigures (a), (b)). The evaluation results of FedNGDB-GQE model facing in-graph queries on FB15k-237 (subfigures (c), (d)).
@@ 5.2. Performance Evaluation
We evaluate FedNGDB’s complex query answering performance on three datasets and compare it to other baselines. We apply FedNGDB on three base query encoding models and evaluate the average retrieval performance on various queries. The results are summarized in Table 3 and Table 4. Table 3 reports the retrieval performance of various distributed graph complex query answering models when there are 3 clients. For each model, we evaluate performance facing in-graph queries and cross-graph queries, respectively. For in-graph queries, the average scores of all clients are reported. We report results in HR@3 and MRR which higher scores indicate better performance. The best results are underlined. The best results of distributed models are in bold. As shown in Table 3, our proposed methods can effectively retrieve complex query answers from distributed graph databases. In comparison to local settings, we can see that FedNGDB can utilize all participated local graph databases and performs better facing in-graph queries. For example, GQE model with FedNGDB can achieve 14.21 HR@3 on average while can only reach 11.85 without collaboration. Besides, in comparison to other federated knowledge graph embedding methods, our proposed FedNGDB can reach comparable performance in both in-graph queries and cross-graph queries without exposing sensitive entity embeddings to the server. For example, FedNGDB achieves the best performance in cross-graph queries in more than half of datasets and base query encoding models.
In Table 4, we present the performance of FedNGDB when there are 5 clients. We compare the performance with local training without collaboration to demonstrate the influence of client numbers. As shown in the table, FedNGDB performs well compared to complex query answering models without collaboration, there are performance improvements in all datasets after applying FedNGDB to various base query encoding models, demonstrating that FedNGDB can utilize the intrinsic information in the distributed knowledge sets. The collaboration allows FedNGDB to reason over various logical paths to improve performance.
@@ 5.3. Query Types
FedNGDB can globally reason over distributed graphs and retrieve answers to cross-graph queries. To evaluate FedNGDB’s performance on various types of complex queries, we conduct experiments to evaluate the retrieval performance of FedNGDB and compare it to central learning on FB15k-237 when there are 3 clients. Figure 4(a) and Figure 4(b) show the FedNGDB-GQE’s performance facing cross-graph queries. As we can see from the figure, FedNGDB performs well on most various types of queries compared to the central model. For example, on query types ’2i’ and ’pi’, FedNGDB can reach more than 90% MRR compared to the central model.
@@ 5.4. Local Influence
Because the in-graph queries can be processed by a single local neural graph database, in this part, we evaluate the performance of FedNGDB on these queries to assess the influence of FedNGDB on local queries. We conduct experiments on FB15k-237 and the number of clients is 5. We evaluate the performance of FedNGDB based on GQE and compare the model with no collaboration. The results are summarized in the Figure 4(c) and Figure 4(d). As shown in the figure, although each client has different performance due to the triplets correlation being different in each sub-dataset, FedNGDB can improve all clients’ performance compared to the model without collaboration.
5.5. Convergence Rate
We evaluate the convergence speed of three federated frameworks. The results are presented by the average number of communication round ratios relative to FedE. As shown in Table 5, FedC’s convergence speed is faster than FedC while slightly slower than FedE, demonstrating that our FedNGDB can protect stronger protection while remaining competitive efficiency.
Table 5. The statistics of communication rounds .
Setting FedE FedR FedNGDB
Relative Rounds to FedE 1.00 1.32 1.09
@@ 6. Conclusion
In this work, we present a federated neural graph database, FedNGDB to reason over distributed knowledge sets with privacy preservation, allowing graph database holders to collaboratively build a distributed graph reasoning system without sharing raw data. We define the distributed graph complex logical query answering problem. To solve the problem, we propose secret aggregation for federated learning where the aggregated parameters can be kept secret to the server. Besides, we design a distributed query retrieval process for answering queries from distributed graph database sets to protect clients’ privacy. To evaluate FedNGDB model performance, we construct a benchmark based on three commonly used knowledge graph complex query answering datasets: FB15k-237, FB15k, and NELL995. Extensive experiments on the benchmark demonstrate the effectiveness of our proposed FedNGDB. FedNGDB can retrieve answers given a query while keeping the sensitive information secret at local graph databases. In the future, we aim to propose new methods for better answering complex queries by exploiting intrinsic information in the distributed neural graph databases.
@@ Appendix A Parameter Sharing
  • Client B chooses a secret integer $b$, then sends Client A $m_{B}=g^{b}\mod p$.
@@
Appendix B Alogrithm
B.1. Secret Aggregation
We present the pseudo-code of secret aggregation in Algorithm 1.
10:  $n$ clients $C_{1},C_{2},\ldots,C_{n}$, client $C_{i}$ has parameter $\theta_{i}$
  Each client $C_{i}$ generates random parameter $\theta_{i}^{r}$
2  Transmit $\theta_{i}^{r}$ to all other clients with encryption, each
3client has a set of parameters $\{\theta_{1}^{r},\theta_{2}^{r}.\cdots,\theta_{n}^{r}\}$
  Client $C_{i}$:
  Upload perturbed parameter $(\theta_{i}+\theta_{i}^{r})$ to server with HE Encryption.
  Receive encrypted parameters $\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})$ from server and HE decryption.
4  Compute encrypted parameters $\theta=[\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})-\sum_{i}^{n}\theta_{i}^{r}]/n$
  Server:
  for $i=1,\cdots,n$ do
     Receive encrypted parameters $(\theta_{i}+\theta_{i}^{r})$ from client $C_{i}$
  end for
  Compute encrypted parameters $\theta^{r}=\sum_{j=1}^{n}(\theta_{j}+\theta_{j}^{r})$
  Send encrypted parameters $\theta^{r}$ to all clients
Algorithm 1 Secret Aggregation
B.2. FedNGDB Framework
We present the pseudo-code of FedNGDB in Algorithm 2.
Appendix C Auxiliary Experiments
Here we present some auxiliary experiments to further evaluate the performance of FedNGDB.
C.1. More Clients
In the former experiments, we evaluate FedNGDB’s performance when there are 3 or 5 clients in the federated system. To further evaluate models’ performance when there are more clients. We split the graph-structured data into 10 subgraphs and evaluate the retrieval performance of FedNGDB using GQE as the base model. As shown in Table 6, FedNGDB can still improve the retrieval performance compared to local training when there are more clients participating in the distributed system, demonstrating the effectiveness of FedNGDB.
@@ C.1. More Clients
2selected in each round $F$;
@@
3  Upload $\mathbf{E}^{i}$ to server for secret aggregation, receive $\mathbf{E}_{0}^{i}$
@@
Algorithm 2 FedNGDB Framework
Table 6. The performance (MRR) of GQE when #C=10.
Setting FB15k-237 FB15k NELL995
Local 8.46 13.04 7.83
FedNGDB 10.17 14.98 9.17
C.2. Relation Overlap
In the experiments, we evaluate the retrieval performance when there are no overlap relations between local graph databases, however, various graph databases can have shared relations in real life, to evaluate the performance in such a scenario, we evaluate the FedNGDB with GQE’s retrieval performance. The graph-structured data is randomly split into 3 subgraphs. The results are shown in Table 7, showing that FedNGDB can successfully retrieve answers from distributed graph databases.
Table 7. The MRR of GQE when relation overlapped.
Setting FB15k-237 FB15k NELL995
Local 10.22 20.21 9.64
FedNGDB 11.42 22.47 11.36